OpenStudy (anonymous):

what value of x makes this proportion true 120/165=x/11

OpenStudy (anonymous):

make it a proportion!!! :D

OpenStudy (anonymous):

120*11 which is 1320 and divide that by 165 and you get 8 so 8 is the correct answer

OpenStudy (anonymous):

wait let me see

OpenStudy (triciaal):

165/11 = so you had to divide 165 by this number to get 11 divide 120 by this number = x

OpenStudy (anonymous):

@triciaal i just said that right?

OpenStudy (anonymous):

120/165=0.7272 x/11=0.7272 0.7272*11=x x=8 (rounded)

OpenStudy (anonymous):

lmaoo thx everyone

OpenStudy (anonymous):

@arzunoorali is right!